OpenStudy (anonymous):

State the x-coordinate of the x-intercept in quadrant 1 for the function f(x)=x^3-9x

OpenStudy (anonymous):

you are being asked to solve \[x^3-9x=0\] or \[x(x^2-9)=0\] i.e. \[x(x+3)(x-3)=0\]